A Basic HyperTEX Example

This is a very basic HyperTEX document. Although DVI files are inherently linear in organization, basic linking can greatly expand their functionality, expecially when used in conjuction with a web browser.

It goes without saying, the math turns out nicely laid out:

∑_{i=0}^{∞} \frac{1}{n^{d}} = ∏_{primes p} (1 + \frac{1}{p^{d}} )

Illustrated below are a variety of link types:

• Links within DVI files

• ⊤Plain HTML files

• ⊤Animations

• ⊤Java pages This is a second page. It exists only to provide a location for a sample anchor.

2